# Regularized Online RLHF with Generalized Bilinear Preferences

### v1

@@ D.1 Four Regret Definitions and Discussions
The four regrets are defined as follows. (a) Average-Nash Regret: $\mathrm{AN\text{-}Reg}_{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\eta}(T):=\max_{\pi^{1},\pi^{2}\in\Pi}\sum_{t=1}^{T}\left\{J_{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\eta}(\pi^{1},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\hat{\pi}_{t}^{2}})-J_{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\eta}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\hat{\pi}_{t}^{1}},\pi^{2})\right\}.$ (b) Average-Best-Response Regret: $\mathrm{ABR\text{-}Reg}_{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\eta}(T):=\sum_{t=1}^{T}\max_{\pi^{1},\pi^{2}\in\Pi}\left\{J_{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\eta}(\pi^{1},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\hat{\pi}_{t}^{2}})-J_{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\eta}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\hat{\pi}_{t}^{1}},\pi^{2})\right\}.$ (c) Max-Nash Regret: $\mathrm{MN\text{-}Reg}_{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\eta}(T):=\max_{\pi\in\Pi}\sum_{t=1}^{T}\left\{\frac{1}{2}-J_{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\eta}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\hat{\pi}_{t}^{1}},\pi)\right\}.$ (d) Max-Best-Response Regret: $\mathrm{MBR\text{-}Reg}_{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\eta}(T):=\sum_{t=1}^{T}\max_{\pi\in\Pi}\left\{\frac{1}{2}-J_{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\eta}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\hat{\pi}_{t}^{1}},\pi)\right\}.$ The unregularized variants are defined similarly and denoted without the $\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\eta$} subscript.
@@ Appendix E Eluder Dimension of GBPM
For GBPM, the eluder dimension $d(\mathcal{P},\lambda,T)$ is bounded as $d(\mathcal{P},\lambda,T)\leq\frac{2d^{2}L_{\mu}^{2}}{\kappa^{2}}\log\left(1+\frac{4\kappa^{2}S^{2}T}{d^{2}\lambda}\right).$
@@ E.2 Connection to the Standard Eluder Dimensions
Suppose that ${\mathcal{Z}}\subseteq{\mathcal{B}}^{d}(1)$ and $\Theta\subseteq\mathrm{Skew}(d;2r,S)$ for some $S>0$. Let $\varepsilon>0$, $T\geq\mathrm{Edim}({\mathcal{P}},\varepsilon)$, and $\lambda\geq 1$. Then, $\frac{\varepsilon^{2}\mathrm{Edim}({\mathcal{P}},\varepsilon)}{\lambda+\varepsilon^{2}}\leq\mathrm{SEC}_{\lambda}({\mathcal{P}},{\bm{D}},T)\lesssim\mathrm{Edim}({\mathcal{P}},T^{-1/2})\log T.$ (38)
@@
Let ${\mathcal{X}}\subseteq{\mathcal{B}}^{d}(1)$ and $\Theta\subseteq\mathrm{Skew}(d,2r;S)$ for some $S>0$. Define the function class as the GBPM: ${\mathcal{P}}:=\left\{({\bm{x}},{\bm{y}})\mapsto\mu\left({\bm{x}}^{\top}\bm{\Theta}{\bm{y}}\right):\bm{\Theta}\in\Theta\right\},$ (39) where we have the following properties: $\kappa\leq\dot{\mu}({\bm{x}}^{\top}\bm{\Theta}{\bm{y}})\leq L_{\mu}$ and $\left|{\bm{x}}^{\top}\bm{\Theta}{\bm{y}}\right|\leq S$ [^11] This follows from matrix Hölder inequality: $\left|{\bm{x}}^{\top}\bm{\Theta}{\bm{y}}\right|=\left|\langle\bm{\Theta},{\bm{x}}{\bm{y}}^{\top}\rangle\right|\leq\left\lVert\bm{\Theta}\right\rVert_{\mathrm{nuc}}\left\lVert{\bm{x}}{\bm{y}}^{\top}\right\rVert_{\mathrm{op}}\leq S$. for all $({\bm{x}},{\bm{y}},\bm{\Theta})\in{\mathcal{X}}\times{\mathcal{X}}\times\Theta$. Then, its eluder dimension is bounded as follows: for any $\varepsilon<SL_{\mu}$, $\mathrm{Edim}({\mathcal{P}},\varepsilon)\leq\frac{3e}{e-1}\cdot d^{2}\cdot\frac{L_{\mu}^{2}}{\kappa^{2}}\cdot\log\frac{24S^{2}L_{\mu}^{2}}{\varepsilon^{2}}.$ (40) Furthermore, if $\mu(z)=\frac{1}{2}+z$, then we have the following lower bound: $\mathrm{Edim}({\mathcal{P}},\varepsilon)\geq\binom{d}{2}\log_{4}\frac{S}{\sqrt{3}\varepsilon}.$ (41)
@@ F.1 Regret Bound of Greedy Sampling
For any $\delta\in(0,1)$, with probability with at least $1-\delta$, GS (when applied to GBPM) attains the following regret bound: $\mathrm{MBR\text{-}Reg}_{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\eta}(T)\lesssim{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}e^{\eta}}\cdot\frac{d^{2}L_{\mu}^{2}}{\kappa^{2}}\cdot\log T\cdot\left(\log\frac{T}{\delta}+dr\log(LST)\right).$
@@
Suppose $D_{\mathrm{ref}}:=\max_{\pi\in\Pi}D_{\mathrm{KL}}(\pi,\pi_{\mathrm{ref}})<\infty$. Then with probability at least $1-\delta$, the unregularized regret of GS under GBPM satisfies $\mathrm{MBR\text{-}Reg}(T)\lesssim{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}e^{\eta}}d^{2}\kappa^{-2}L_{\mu}^{2}\cdot\log T\cdot\left(\log\frac{T}{\delta}+dr\log(LST)\right)+{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\eta^{-1}}D_{\mathrm{ref}}T.$ Furthermore, there exists no $\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\eta$} (even dependent on $T$) such that the RHS is ${\mathcal{O}}(T^{1-\gamma})$ for any $\gamma\in(0,1]$.

### v2

@@ D.1 Four Regret Definitions and Discussions
The four regrets are defined as follows.
  1. (a)
    Average-Nash Regret:
    $\mathrm{AN\text{-}Reg}_{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\eta}(T):=\max_{\pi^{1},\pi^{2}\in\Pi}\sum_{t=1}^{T}\left\{J_{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\eta}(\pi^{1},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\hat{\pi}_{t}^{2}})-J_{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\eta}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\hat{\pi}_{t}^{1}},\pi^{2})\right\}.$
  2. (b)
    Average-Best-Response Regret:
    $\mathrm{ABR\text{-}Reg}_{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\eta}(T):=\sum_{t=1}^{T}\max_{\pi^{1},\pi^{2}\in\Pi}\left\{J_{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\eta}(\pi^{1},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\hat{\pi}_{t}^{2}})-J_{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\eta}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\hat{\pi}_{t}^{1}},\pi^{2})\right\}.$
  3. (c)
    Max-Nash Regret:
    $\mathrm{MN\text{-}Reg}_{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\eta}(T):=\max_{\pi\in\Pi}\sum_{t=1}^{T}\left\{\frac{1}{2}-J_{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\eta}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\hat{\pi}_{t}^{1}},\pi)\right\}.$
  4. (d)
    Max-Best-Response Regret:
    $\mathrm{MBR\text{-}Reg}_{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\eta}(T):=\sum_{t=1}^{T}\max_{\pi\in\Pi}\left\{\frac{1}{2}-J_{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\eta}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\hat{\pi}_{t}^{1}},\pi)\right\}.$
The unregularized variants are defined similarly and denoted without the $\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\eta$} subscript.
@@ Appendix E Eluder Dimension of GBPM
For GBPM, the eluder dimension $d(\mathcal{P},\lambda,T)$ is bounded as
$d(\mathcal{P},\lambda,T)\leq\frac{2d^{2}L_{\mu}^{2}}{\kappa^{2}}\log\left(1+\frac{4\kappa^{2}S^{2}T}{d^{2}\lambda}\right).$
@@ E.2 Connection to the Standard Eluder Dimensions
Suppose that ${\mathcal{Z}}\subseteq{\mathcal{B}}^{d}(1)$ and $\Theta\subseteq\mathrm{Skew}(d;2r,S)$ for some $S>0$. Let $\varepsilon>0$, $T\geq\mathrm{Edim}({\mathcal{P}},\varepsilon)$, and $\lambda\geq 1$. Then,
$\frac{\varepsilon^{2}\mathrm{Edim}({\mathcal{P}},\varepsilon)}{\lambda+\varepsilon^{2}}\leq\mathrm{SEC}_{\lambda}({\mathcal{P}},{\bm{D}},T)\lesssim\mathrm{Edim}({\mathcal{P}},T^{-1/2})\log T.$ (38)
@@
Let ${\mathcal{X}}\subseteq{\mathcal{B}}^{d}(1)$ and $\Theta\subseteq\mathrm{Skew}(d,2r;S)$ for some $S>0$. Define the function class as the GBPM:
${\mathcal{P}}:=\left\{({\bm{x}},{\bm{y}})\mapsto\mu\left({\bm{x}}^{\top}\bm{\Theta}{\bm{y}}\right):\bm{\Theta}\in\Theta\right\},$ (39)
where we have the following properties: $\kappa\leq\dot{\mu}({\bm{x}}^{\top}\bm{\Theta}{\bm{y}})\leq L_{\mu}$ and $\left|{\bm{x}}^{\top}\bm{\Theta}{\bm{y}}\right|\leq S$ [^11] This follows from matrix Hölder inequality: $\left|{\bm{x}}^{\top}\bm{\Theta}{\bm{y}}\right|=\left|\langle\bm{\Theta},{\bm{x}}{\bm{y}}^{\top}\rangle\right|\leq\left\lVert\bm{\Theta}\right\rVert_{\mathrm{nuc}}\left\lVert{\bm{x}}{\bm{y}}^{\top}\right\rVert_{\mathrm{op}}\leq S$. for all $({\bm{x}},{\bm{y}},\bm{\Theta})\in{\mathcal{X}}\times{\mathcal{X}}\times\Theta$. Then, its eluder dimension is bounded as follows: for any $\varepsilon<SL_{\mu}$,
$\mathrm{Edim}({\mathcal{P}},\varepsilon)\leq\frac{3e}{e-1}\cdot d^{2}\cdot\frac{L_{\mu}^{2}}{\kappa^{2}}\cdot\log\frac{24S^{2}L_{\mu}^{2}}{\varepsilon^{2}}.$ (40)
Furthermore, if $\mu(z)=\frac{1}{2}+z$, then we have the following lower bound:
$\mathrm{Edim}({\mathcal{P}},\varepsilon)\geq\binom{d}{2}\log_{4}\frac{S}{\sqrt{3}\varepsilon}.$ (41)
@@ F.1 Regret Bound of Greedy Sampling
For any $\delta\in(0,1)$, with probability with at least $1-\delta$, GS (when applied to GBPM) attains the following regret bound:
$\mathrm{MBR\text{-}Reg}_{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\eta}(T)\lesssim{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}e^{\eta}}\cdot\frac{d^{2}L_{\mu}^{2}}{\kappa^{2}}\cdot\log T\cdot\left(\log\frac{T}{\delta}+dr\log(LST)\right).$
@@
Suppose $D_{\mathrm{ref}}:=\max_{\pi\in\Pi}D_{\mathrm{KL}}(\pi,\pi_{\mathrm{ref}})<\infty$. Then with probability at least $1-\delta$, the unregularized regret of GS under GBPM satisfies
$\mathrm{MBR\text{-}Reg}(T)\lesssim{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}e^{\eta}}d^{2}\kappa^{-2}L_{\mu}^{2}\cdot\log T\cdot\left(\log\frac{T}{\delta}+dr\log(LST)\right)+{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\eta^{-1}}D_{\mathrm{ref}}T.$
Furthermore, there exists no $\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\eta$} (even dependent on $T$) such that the RHS is ${\mathcal{O}}(T^{1-\gamma})$ for any $\gamma\in(0,1]$.